# Final-Model-Only Data Attribution with a Unifying View of Gradient-Based Methods

### v1

@@ 6.2 Results
Figure 2: Cosine similarity between attribution scores of gradient-based TDA methods and further training, as a function of the number of epochs or steps of further training. The legend in panel (b) applies to (a)–(d). Note that in (c), CG overlaps almost perfectly with LiSSA.
@@
Figure 2 shows the cosine similarity between the attribution scores of gradient-based TDA methods and further training, as a function of the amount of further training. The curves represent the mean cosine similarity over the $m$ test instances, while the shaded area corresponds to one standard error above and below.
We make the following observations:
  • First-order methods: The two first-order methods, Grad-Dot and Grad-Cos, have cosine similarity curves that decay with the amount of further training. Grad-Dot is generally superior to Grad-Cos, and the initial cosine similarity of the former is among the highest achieved by any approximate method. This aligns with Section 4.2 showing that Grad-Dot is the direct result of first-order approximation of further training, whereas the normalization of Grad-Cos is not supported by the theory. The notable exception is for CIFAR-10 (Figure 2(e)), which we comment on in Appendix E. The rate of decay after the initial peak varies with the dataset. One reason may be how far from stationary are the final model parameters ${\bm{\theta}}^{f}$. If they are farther, then the parameter changes $\Delta{\bm{\theta}}(\mathcal{D})$ and $\Delta{\bm{\theta}}(\mathcal{D}_{-i})$ are larger, the first-order approximation is poorer, and the cosine similarity is thus lower.
  • DataInf: We find it interesting that DataInf behaves like a first-order method, despite its attempt to incorporate second-order information. It is especially similar to Grad-Dot (cosine similarity between the two is typically above $0.95$) and generally slightly higher in similarity to further training. We further interpret this similarity between DataInf and Grad-Dot in Appendix E.
@@
  • EK-FAC, TRAKM=1: These are influence function methods that make additional approximations. Compared to LiSSA, EK-FAC has similar or higher cosine similarity in the second row of Figure 2, but is mostly worse in the first row. TRAKM=1 did not give good approximations to further training in any of our experiments.
  • CIFAR-10, SST-2: The cosine similarities are significantly lower for these non-tabular datasets and models. The immediate reason for this is that the gold attribution vectors ${\bm{a}}$ behave differently than they do for tabular data, as discussed in Appendix E.
(a) Concrete Strength
(b) Energy Efficiency
(c) FICO
(d) Folktables
(e) CIFAR-10
(f) SST-2
@@
Similarity vs. amount of averaging
In Figure 3, we plot the maximum cosine similarity over epochs/steps (i.e., the maximum of each curve in Figure 2) as a function of the number $r$ of random seeds averaged. The procedure is as follows. Given a number $r$ of seeds to be averaged, we divide the 100 seeds that we have in total into $100/r$ groups of $r$ seeds. For each group of $r$ seeds, we obtain a gold attribution vector ${\bm{a}}$ using (13) and compute its cosine similarity with a gradient-based attribution vector $\hat{{\bm{a}}}$. We then compute means and standard errors of the cosine similarities over the $m$ test instances as well as the $100/r$ groups. This yields mean cosine similarities as a function of further training epochs or steps, like in Figure 2. Finally, we take the maximum over epochs or steps to give us the $y$-axis value in Figure 3 corresponding to the number of seeds $r$ on the $x$-axis. This procedure is repeated for different values of $r$.
In all cases in Figure 3, the cosine similarity increases with the number of seeds, implying that the averaging of further training runs brings it closer to what gradient-based methods estimate. This supports our proposal of averaging and its theoretical counterpart, the expectation in (3). While in some cases the increases are not visible in the plots, they are more notable for the first-order methods and DataInf, and for certain datasets like SST-2.
7 Conclusion
We have highlighted the final-model-only setting for TDA and proposed a further training gold standard for this setting. We then showed how several gradient-based TDA methods approximate further training in different ways, theoretically and numerically.
We note the following possibilities for future work:
  1. 1)
    Given the cosine similarity results in Figure 2, it is natural to hope for higher-quality approximate TDA methods, especially for non-tabular models. One possible starting point is the generalized influence function expressions in Proposition 1 and Corollary 1, which suggest a potential role for the parameter change $\widehat{\Delta{\bm{\theta}}}(\mathcal{D})$ from further training on the full dataset.
  2. 2)
    Figure 2 also suggests the question of whether the strengths of first- and second-order methods could be combined, i.e., higher initial similarity with less decay. We note that the damping parameter $\lambda$ could be used to interpolate between first- and second-order methods.
  3. 3)
    Another open question is how much further training is the right amount to determine the final model’s sensitivity to training instances. If too little, the sensitivity may not be measurable, whereas if too much, then the measurement may no longer be local, i.e., specific to the given model $f({\bm{x}};{\bm{\theta}}^{f})$. The extent to which further training is well-approximated by Taylor expansions (as evaluated for example in Figure 2) may provide a partial answer.
  4. 4)
    We restricted attention in this work to LOO further training, as mentioned in Section 3. Future research on TDA in general could focus more on units larger than a single instance (referred to as “group influence” (Koh et al., 2019; Basu et al., 2020)), beyond making the additivity assumption that the attribution score for a group is the sum of LOO attribution scores.
  5. 5)
    While we have presented further training as applied to the original training set $\mathcal{D}$ and assumed that $\mathcal{D}$ is available, there is no obstacle to considering further training on new, unseen instances. This would change the problem from TDA to predicting the effect of further training on new data.
@@ Acknowledgements
We thank Swagatam Haldar for early discussions on the TDA literature.
@@ Appendix A Further Training with Convex Objectives
Given the relative ease of obtaining the exact minimizer ${\bm{\theta}}^{*}(\mathcal{D})$ on the full training set, it is reasonable to assume that any final, trained parameters have converged to it, i.e., ${\bm{\theta}}^{f}={\bm{\theta}}^{*}(\mathcal{D})$. In this case, there is no non-convergence problem as in Section 3, and further training would yield no change, i.e., $\Delta{\bm{\theta}}(\mathcal{D})={\bm{0}}$ in (1). As for stochastic training algorithms, they can of course also be applied to convex objectives, often with convergence guarantees (see e.g. Shamir and Zhang (2013) as one example among many). This greatly reduces the problem of stochasticity.
@@ B.1 Regularization/damping term
    Enforce assumption/trust region: Penalizing the $\ell_{2}$ norm of $\Delta{\bm{\theta}}$ enforces the assumption that $\Delta{\bm{\theta}}$ is small. It can also be viewed as enforcing a trust region (Conn et al., 2000), a region around ${\bm{\theta}}^{f}$ where the second-order expansion gives a better approximation.
@@
    Ensure optimization is well-posed: The loss function $L({\bm{z}};{\bm{\theta}})$ for neural networks is typically non-convex, which implies that the empirical risk Hessian $\nabla^{2}_{\bm{\theta}}R(\mathcal{D}^{\prime};{\bm{\theta}}^{f})$ may not be positive semidefinite. In the absence of the regularizer, if $\nabla^{2}_{\bm{\theta}}R(\mathcal{D}^{\prime};{\bm{\theta}}^{f})$ has a negative eigenvalue, then the objective value in (4) would tend to $\infty$- ∞ in the direction of the corresponding eigenvector. If we add regularization with $\lambda$ larger in magnitude than the most negative eigenvalue of $\nabla^{2}_{\bm{\theta}}R(\mathcal{D}^{\prime};{\bm{\theta}}^{f})$, then such divergence is prevented and an optimal solution to (4) exists. This consideration leads to Assumption 1.
@@ Appendix C More on Related Work
Bae et al. (2022) examine the discrepancy between influence functions and leave-one-out re-training and decompose the discrepancy into five contributions. They show that influence functions are a much better approximation to a quantity they call the proximal Bregman response function (PBRF), which they propose as an alternative gold standard to LOO re-training. The key differences between our work and Bae et al. (2022) are as follows: 1) We propose further training based on standard training objectives (e.g., cross-entropy or mean squared error, no proximity regularization) as a different, more general gold standard. In contrast, the PBRF involves a non-standard Bregman distance objective, chosen specifically to be closer to influence functions. In their Appendix D.2, Bae et al. (2022) also discuss further training with and without leaving out one sample, which they refer to as “two-stage LOO re-training.” However, they do not consider taking an expectation over random trainings as we do. 2) Our unified view encompasses TDA methods beyond influence functions, including more recent methods such as TRAK (Park et al., 2023) and DataInf (Kwon et al., 2024) as well as first-order methods. We evaluate the extent to which all these methods approximate further training.
@@
Schioppa et al. (2023) discuss problematic assumptions made by influence functions and TracIn (Pruthi et al., 2020) and how these can be addressed. They then show that the predictive power of influence functions and TracIn is fundamentally limited by divergence in parameters as (further) training proceeds. The latter contribution is the main point of connection with our work, where we also observe the fading of predictive power over training time, especially for first-order methods. The differences in our work are: 1) We make explicit the further training gold standard that they compare to in their experiment (Schioppa et al., 2023, Sec. 5.2) and we further propose its expected version with respect to stochasticity in the training algorithm. 2) We conduct a larger-scale experiment than theirs in certain dimensions, notably the TDA methods evaluated, as well as averaging over the aforementioned stochasticity and using more test points and left-out training points.
@@
Basu et al. (2021) show that influence functions become worse approximators of a training example’s importance on a test prediction with increasing model depth, width and if the models are trained with no weight decay. All experiments in the paper are conducted with further training (from the already trained model, not from scratch) determining the ground truth influence. However, Basu et al. (2021) do not average over further training realizations as we do. Moreover, their study is focused on influence functions, whereas we consider other gradient-based methods as well.
@@
Nguyen et al. (2023) argue that a Bayesian approach is the correct way of evaluating influences of training examples on test predictions as there is a lot of variance in most TDA estimates when we account for random initialization of the model as well as variation in batching when training. In such situations the recommendation is to consider the distribution of attribution values for a training example, or at least consider the distribution’s variance, rather than simply (individual run) point estimates. Different from our work, the experiments in the paper are conducted by re-training models from scratch, and only three gradient-based TDA methods are studied, namely influence functions (IF), Grad-Dot (GD) and Grad-Cos (GC).
@@
The seminal paper (Koh and Liang, 2017) on influence functions in ML sketched a derivation similar to ours in Section 4 and Appendix B, but in less detail. In particular, our more careful derivation of the influence function in Proposition 1 keeps the $\widehat{\Delta{\bm{\theta}}}(\mathcal{D})$ term that Koh and Liang (2017) neglect. Appendix B.1 also discusses the damping term $\lambda{\bm{I}}$ at greater length.
@@
Zhang and Zhang (2022) analyze the effectiveness of influence functions, but only for a 2-layer ReLU network. Moreover, the theoretical analysis is done w.r.t. the neural tangent kernel (NTK) approximation of the ReLU network. They have three key findings: i) Influence functions perform better as the width of the network increases. ii) Influence functions perform poorly with weak regularization, where the regularization tries to keep the model weights as close as possible to the initial weights. Our cosine similarity results in Figure 2 and elsewhere may be a reflection of this. iii) Influence functions perform better for examples that lie in the high density region of the training distribution.
@@ D.1 Data
We followed Bae et al. (2022) in using two regression datasets from the UCI repository (Kelly et al., ), Concrete Compressive Strength and Energy Efficiency. We then departed from Bae et al. (2022) in choosing two classification datasets, FICO Challenge (FICO, 2018) and Folktables (Ding et al., 2021), because they are at least one order of magnitude larger than the tabular datasets in Bae et al. (2022) in terms of the number of instances and/or features.
@@
For Concrete, Energy, and FICO, we split the dataset 90%-10% into training and test sets (using the scikit-learn (Pedregosa et al., 2011) package’s train_test_split() with random_state=0) and standardized the features to have zero mean and unit variance. Handling of special feature values in FICO was done as in Wei (2021). For Folktables, we used the person-level data from year 2018 for the US state of Massachusetts. The classification task was predicting whether a person’s income is above or below USD 50,000 (task ACSIncome). The dataset was split 75%-25% into training and test (again using train_test_split() with random_state=0), categorical features were dummy-coded, and numerical features were standardized. For CIFAR-10, we used the given split into training and test sets. We employed common data augmentation techniques such as random cropping and horizontal flipping, followed by standardization using the empirical mean and standard deviation of the training set. We used a random subset of 1000 samples as validation set. For SST-2, we used the given split into training, validation, and test sets. We drew the $m=100$ “test” instances for evaluation from the validation set because the actual test set lacks labels for computing losses. Text was tokenized using the tokenizer of the BERT model (see next subsection) with a maximum sequence length of 128 (more than sufficient for all SST-2 examples).

### v2

@@ 6.2 Results
Figure 2: Cosine similarity between attribution scores of gradient-based TDA methods and further training, as a function of the number of epochs or steps of further training. The legend in panel (b) applies to (a)–(d). In panel (a), Grad-Cos occludes Grad-Dot; in panel (c), LiSSA occludes CG.
@@
Figure 2 shows the cosine similarity between the attribution scores of gradient-based TDA methods and further training, as a function of the amount of further training. The curves represent the mean cosine similarity over the $m$ test instances, while the shaded area corresponds to $\pm 1$ standard error. We make the following observations:
  • First-order methods: The two first-order methods, Grad-Dot and Grad-Cos, have cosine similarity curves that decay with the amount of further training. Grad-Dot is generally superior to Grad-Cos, and the initial cosine similarity of the former is among the highest of any approximate method. This aligns with Section 4.2 showing that Grad-Dot is the direct result of first-order approximation of further training, whereas the normalization of Grad-Cos is not supported by the theory. The rate of decay after the initial peak varies with the dataset. One reason may be how far from stationary are the final model parameters ${\bm{\theta}}^{f}$. If they are farther, then the parameter changes $\Delta{\bm{\theta}}(\mathcal{D})$ and $\Delta{\bm{\theta}}(\mathcal{D}_{-i})$ are larger, the first-order approximation is poorer, and the cosine similarity is thus lower.
  • DataInf: We find it interesting that DataInf behaves like a first-order method, despite its attempt to incorporate second-order information. It is especially similar to Grad-Dot (cosine similarity between the two is typically above $0.95$) and generally slightly higher in similarity to further training. We further interpret this similarity between DataInf and Grad-Dot in Appendix E.1.
@@
  • EK-FAC, TRAKM=1: These influence function methods make additional approximations. Compared to LiSSA, EK-FAC has similar or higher cosine similarity in Figures 2(d)2(f), but is worse in 2(a)2(c). TRAKM=1 did not give good approximations to further training (please see Appendix E.1).
  • CIFAR-10, SST-2: The cosine similarities are significantly lower for these non-tabular datasets. The immediate reason for this is that the gold attribution vectors $\bm{a}$} behave differently than they do for tabular data, as discussed in Appendix E.1.
Similarity vs. amount of averaging
In Figure 3, we plot the maximum cosine similarity over epochs/steps (i.e., the maximum of each curve in Figure 2) as a function of the number $r$ of random seeds averaged in (13) to obtain gold attribution values. The exact averaging procedure is given in Appendix E.1, along with a full version of Figure 3 with all datasets (Figure 7). We find in all cases that the cosine similarities increase with the number of seeds, implying that the averaging of further training runs brings it closer to what gradient-based methods estimate. This supports our proposal of averaging and its theoretical counterpart, the expectation in (3). The increases are more noticeable for the first-order methods and DataInf, and for certain datasets like SST-2.
7 Discussion and Future Work
(a) Folktables
(b) SST-2
@@ 7 Discussion and Future Work
Summary
We have highlighted the final-model-only setting for TDA, recasting the problem (which we call FiMODA) as one of measuring the model’s sensitivity to training instances rather than their contribution. We accordingly proposed further training, with refinements, as a gold standard method for FiMODA. We showed how several existing gradient-based TDA methods approximate further training, theoretically and numerically. These relationships suggest that gradient-based methods are better suited to FiMODA than to other TDA settings.
Limitations
The further training gold standard as described in this work is computationally expensive (as acknowledged at the beginning of Section 4). Performing further training also limited the scale of the experiments in Section 6 in terms of the models considered, number of training instances left out, etc. The computational expense might be mitigated in two ways: First, Figures 3 and 79 suggest that while averaging does bring further training closer to gradient-based methods, most of the gain is achieved with 10 or 20 seeds compared to 100. So reducing the number of random seeds is a computational trade-off that could be made. Second, while we have implemented further training as full fine-tuning, it is possible to use LoRA fine-tuning or other parameter-efficient methods (especially for LLMs) as another computational trade-off. The LoRA Ensemble method of deng2024efficient shows concretely how both ideas, ensembling and LoRA fine-tuning, can be brought together in the context of TDA. We mention other limitations in Appendix F.
Discussion of experimental results
Perhaps the most salient aspect of our experimental results (Figures 2, 3 in Section 6.2 and variations in Appendix E.1) is the difference between the first-order-like methods (Grad-Dot, Grad-Cos, DataInf) and influence-function-based methods. In particular, the former can achieve significantly higher cosine similarities than the latter when the amount of further training is low. This contrasts with previous work in TDA in the TAA or CPA settings, where influence-function-based methods are generally seen as stronger park2023trak; bae2024unrolled. Our results for TRAK may be especially surprising in this regard, for which we point to two factors: First, in the FiMO setting, we have only $M=1$ checkpoint and can only evaluate the TRAKM=1 variant, whereas park2023trak shows that using an ensemble of $M\gg 1$ checkpoints greatly improves performance. Second, we compute similarity with further training attribution values, whereas park2023trak evaluate using their linear datamodelling score (LDS) metric, which is based on re-training and for the TAA setting.
Overall however, we think that our experimental results leave something to be desired for researchers in TDA. It could be argued that the gradient-based methods are “actually similar” to further training only for the first-order-like methods, under limited further training, and perhaps not at all in Figures 2(e), 2(f). (It is perhaps worth recalling that all of these gradient-based methods are existing and we are not advocating for any particular method.) Our results therefore motivate development of higher-quality approximate FiMODA methods, especially for non-tabular models. One possible starting point is the generalized influence function expressions in Proposition 1 and Corollary 2, which suggest a potential role for the parameter change $\widehat{\Delta{\bm{\theta}}}(\mathcal{D})$ from further training on the full dataset. Moreover, Figure 2 suggests the question of whether the strengths of first- and second-order methods could be combined, i.e., higher initial similarity with less decay. We note that the damping parameter $\lambda$ could be used to interpolate between first- and second-order methods.
Additional future work
Appendix F elaborates on the following points: 1) What is the right amount of further training to determine sensitivity to training instances? 2) Section 3 restricted attention to LOO further training to align with gradient-based methods; group influence (koh2019group; basu2020second; hu2024most) is worthy of further study. 3) Further training could also be applied to new, unseen data.
@@ Acknowledgements
We thank Swagatam Haldar for early discussions on the TDA literature. We also thank the anonymous NeurIPS reviewers for their constructive comments and discussions, in particular for clarifying that we are indeed reframing the problem for the FiMO setting and for suggesting an expanded discussion section.
@@ Appendix A Further Training with Convex Objectives
Given the relative ease of obtaining the exact minimizer ${\bm{\theta}}^{*}(\mathcal{D})$ on the full training set, it is reasonable to assume that any final, trained parameters have converged to it, i.e., ${\bm{\theta}}^{f}={\bm{\theta}}^{*}(\mathcal{D})$. In this case, there is no non-convergence problem as in Section 3, and further training would yield no change, i.e., $\Delta{\bm{\theta}}(\mathcal{D})={\bm{0}}$ in (1). As for stochastic training algorithms, they can of course also be applied to convex objectives, often with convergence guarantees (see e.g. [shamir2013sgd] as one example among many). This greatly reduces the problem of stochasticity.
@@ B.1 Regularization/damping term
    Enforce assumption/trust region: Penalizing the $\ell_{2}$ norm of $\Delta{\bm{\theta}}$ enforces the assumption that $\Delta{\bm{\theta}}$ is small. It can also be viewed as enforcing a trust region [conn2000trustregion], a region around ${\bm{\theta}}^{f}$ where the second-order expansion gives a better approximation.
@@
    Ensure optimization is well-posed: The loss function $L({\bm{z}};{\bm{\theta}})$ for neural networks is typically non-convex, which implies that the empirical risk Hessian $\nabla^{2}_{\bm{\theta}}R(\mathcal{D}^{\prime};{\bm{\theta}}^{f})$ may not be positive semidefinite. In the absence of the regularizer, if $\nabla^{2}_{\bm{\theta}}R(\mathcal{D}^{\prime};{\bm{\theta}}^{f})$ has a negative eigenvalue, then the objective value in (4) would tend to -$\infty$ in the direction of the corresponding eigenvector. If we add regularization with $\lambda$ larger in magnitude than the most negative eigenvalue of $\nabla^{2}_{\bm{\theta}}R(\mathcal{D}^{\prime};{\bm{\theta}}^{f})$, then such divergence is prevented and an optimal solution to (4) exists. This consideration leads to Assumption 1.
@@ Appendix C More on Related Work
Bae et al. (2022) [bae2022infanswer] examine the discrepancy between influence functions and leave-one-out re-training and decompose the discrepancy into five contributions. They show that influence functions are a much better approximation to a quantity they call the proximal Bregman response function (PBRF), which they propose as an alternative gold standard to LOO re-training. The key differences between our work and [bae2022infanswer] are as follows: 1) We propose further training based on standard training objectives (e.g., cross-entropy or mean squared error, no proximity regularization) as a different, more general gold standard. In contrast, the PBRF involves a non-standard Bregman distance objective, chosen specifically to be closer to influence functions. In their Appendix D.2, [bae2022infanswer] also discuss further training with and without leaving out one sample, which they refer to as “two-stage LOO re-training.” However, they do not consider taking an expectation over random trainings as we do. 2) Our unified view encompasses TDA methods beyond influence functions, including more recent methods such as TRAK [park2023trak] and DataInf [kwon2024datainf] as well as first-order methods. We evaluate the extent to which all these methods approximate further training. [bae2022infanswer] evaluated two TDA methods, LiSSA with and without the Gauss-Newton Hessian approximation.
@@
Schioppa et al. (2023) [schioppa2023theoretical] discuss problematic assumptions made by influence functions and TracIn [pruthi2020tracin] and how these can be addressed. They then show that the predictive power of influence functions and TracIn is fundamentally limited by divergence in parameters as (further) training proceeds. The latter contribution is the main point of connection with our work, where we also observe the fading of predictive power over training time, especially for first-order methods. The differences in our work are: 1) We make explicit the further training gold standard that they compare to in their experiment [schioppa2023theoretical, Sec. 5.2] and we further propose its expected version with respect to stochasticity in the training algorithm. 2) We conduct a larger-scale experiment than theirs in certain dimensions, notably the TDA methods evaluated, as well as averaging over the aforementioned stochasticity and using more test points and left-out training points. [schioppa2023theoretical] evaluated two methods, TracIn and a Conjugate Residual method of computing influence functions.
@@
Basu et al. (2021) [basu2021inffragile] show that influence functions become worse approximators of a training example’s importance on a test prediction with increasing model depth, width and if the models are trained with no weight decay. All experiments in the paper are conducted with further training (from the already trained model, not from scratch) determining the ground truth influence. This was done for efficiency reasons however, not because of an explicit restriction to the FiMO setting. Moreover, [basu2021inffragile] do not average over further training realizations or correct for non-converged models as we do. Their study is focused on influence functions (computed using two methods, exact and LiSSA), whereas we consider other gradient-based methods as well.
@@
Nguyen et al. (2023) [nguyen2023bayesian] argue that a Bayesian approach is the correct way of evaluating influences of training examples on test predictions as there is a lot of variance in most TDA estimates when we account for random initialization of the model as well as variation in batching when training. In such situations the recommendation is to consider the distribution of attribution values for a training example, or at least consider the distribution’s variance, rather than simply (individual run) point estimates. Different from our work, the experiments in the paper are conducted by re-training models from scratch, and only three gradient-based TDA methods are studied, namely influence functions (IF), Grad-Dot (GD) and Grad-Cos (GC).
@@
The seminal paper by Koh et al. (2017) [koh2017inffunc] on influence functions in ML sketched a derivation similar to ours in Section 4 and Appendix B, but in less detail (in their Section 4.2 and Appendix B, less than half a page in total). In particular, our more careful derivation of the influence function in Proposition 1 keeps the $\widehat{\Delta{\bm{\theta}}}(\mathcal{D})$ term that [koh2017inffunc] neglect. Appendix B.1 also discusses the damping term $\lambda{\bm{I}}$ at greater length. [koh2017inffunc] evaluated two influence function methods, CG and LiSSA (without the Gauss-Newton approximation).
@@
Zhang & Zhang (2022) [zhang2022rethinking] analyze the effectiveness of influence functions, but only for a 2-layer ReLU network. Moreover, the theoretical analysis is done w.r.t. the neural tangent kernel (NTK) approximation of the ReLU network. They have three key findings: i) Influence functions perform better as the width of the network increases. ii) Influence functions perform poorly with weak regularization, where the regularization tries to keep the model weights as close as possible to the initial weights. Our cosine similarity results in Figure 2 and elsewhere may be a reflection of this. iii) Influence functions perform better for examples that lie in the high density region of the training distribution. While the work of [zhang2022rethinking] was mostly theoretical, they did numerically evaluate influence functions computed in two ways, using inverse Hessian-vector products and a simpler method that takes advantage of their NTK approximation.
In addition, we mention Arnoldi influence functions [schioppa2022scalingarnoldi] as a method falling under Section 4.3. [schioppa2022scalingarnoldi] proposed to use Arnoldi iteration (a.k.a. Lanczos iteration for symmetric matrices) to approximately compute the largest-magnitude eigenvalues and corresponding eigenvectors of ${\bm{H}}({\bm{\theta}}^{f})$. The method then projects gradient vectors into the subspace of top eigenvectors of ${\bm{H}}({\bm{\theta}}^{f})$. By virtue of the orthonormal basis of eigenvectors for this subspace, the matrix inversion in (8) reduces to scalar division by eigenvalues.
[…]
@@ D.1 Data
We followed [bae2022infanswer] in using two regression datasets from the UCI repository [uci] (CC BY 4.0 license), Concrete Compressive Strength and Energy Efficiency. We then departed from [bae2022infanswer] in choosing two classification datasets, FICO Challenge [fico2018explainable] (license unknown) and Folktables [ding2021retiring] (MIT license), because they are at least one order of magnitude larger than the tabular datasets in [bae2022infanswer] in terms of the number of instances and/or features.
@@
For Concrete, Energy, and FICO, we split the dataset 90%-10% into training and test sets (using the scikit-learn [scikit-learn] package’s train_test_split() with random_state=0) and standardized the features to have zero mean and unit variance. Handling of special feature values in FICO was done as in [wei2021selective]. For Folktables, we used the person-level data from year 2018 for the US state of Massachusetts. The classification task was predicting whether a person’s income is above or below USD 50,000 (task ACSIncome). The dataset was split 75%-25% into training and test (again using train_test_split() with random_state=0), categorical features were dummy-coded, and numerical features were standardized. For CIFAR-10 (original license unclear), we used the given split into training and test sets. We employed common data augmentation techniques such as random cropping and horizontal flipping, followed by standardization using the empirical mean and standard deviation of the training set. We used a random subset of 1000 samples as validation set. For SST-2 (original license unclear), we used the given split into training, validation, and test sets. We drew the $m=100$ “test” instances for evaluation from the validation set because the actual test set lacks labels for computing losses. Text was tokenized using the tokenizer of the BERT model (see next subsection) with a maximum sequence length of 128 (more than sufficient for all SST-2 examples).